OpenStudy (anonymous):

i am trying to find the area in yds of a triangle with a base of 1 4/5 and height of 5/8 is a= (1/2) (1 4/5)(5/8) a=(1/2)(9/5)(5/8) a=45/80 right and where do i go from there and how

OpenStudy (zzr0ck3r):

area of triangle = 1/2 * b * h you have b = 9/5 and h = 5/8 can you solve?

OpenStudy (anonymous):

Thats where i am stuck it does not make sense to me

OpenStudy (zzr0ck3r):

area = (1/2)*b*h b = 9/5 h = 5/8

OpenStudy (zzr0ck3r):

so area = (1/2)*(9/5)*(5/8) = (1*9*5)/(2*5*8) = 45/?

OpenStudy (anonymous):

80

OpenStudy (anonymous):

Where did you get these values from? a= (1/2) (1 4/5)(5/8) a=(1/2)(9/5)(5/8) a=45/80

OpenStudy (anonymous):

are they wrong

OpenStudy (anonymous):

Ignore me. I see what you did now... yes, the answer is 45/80 = .5625 = 9/16

OpenStudy (anonymous):

thank you i was forgetting to simplify

OpenStudy (anonymous):

No problem! And thank you! :-)